# Convergence for Natural Policy Gradient on Infinite-State Average-Reward Markov Decision Processes

### v1

@@ 4. Results
We prove the first convergence result for the Natural Policy Gradient algorithm in an MDP setting with infinite state space and average-reward objective.
Theorem 1 ().
@@
Proof deferred to Section 5.
To give a concrete example of an infinite-state average-reward MDP satisfying Assumptions 1 and 2 and an initial policy satisfying (1), we examine the “Generalized Switch with Static Environment” (GSSE) in Section 6. This is a highly general queueing model which can capture a wide variety of commonly studied queueing systems, including the N-system, the switch, and the multiserver-job system. We study the MaxWeight initial policy in this setting.
In Theorem 3, we prove that in the GSSE setting, with the MaxWeight initial policy, the NPG algorithm converges to the policy with optimal mean queue length with convergence rate $O(1/\sqrt{T})$.
@@ 4.1. Challenges
To prove a convergence rate result for the NPG algorithm in the infinite-state-space, average-reward, setting, a natural approach would be to try to generalize finite-state average-reward results to the infinite-reward setting. There are two relevant approaches to consider: the approach of Even-Dar et al. (2009), and of Murthy and Srikant (2023). Unfortunately, each of these approaches makes crucial use of assumptions that are only plausible for MDPs with finite state-spaces.
With respect to Even-Dar et al. (2009), combining that paper’s Algorithm 4 (MDP Experts) with its Algorithm 1 (Weighted Majority) results in exactly the Natural Policy Gradient algorithm, and the paper’s Theorem 4.1 proves that NPG converges to the optimal policy in the finite-state average-reward setting, with convergence rate $O(1/\sqrt{T})$. To prove this theorem, the paper assumes that there exists some finite mixing time $\tau$ for all policies $\pi$. Intuitively, this assumption states that every policy takes the system state to a distribution becomes near the policy’s stationary distribution in at most $O(\tau)$ steps. In a finite-state MDP, this is a reasonable assumption: For reasonable policies, $\tau$ might be exponential in $|S|$ at worst. In contrast, in an infinite-state MDP with infinite diameter, such as the queueing systems we consider in Section 6, it is plainly false. No policy can achieve a finite mixing time, much less all policies.
Murthy and Srikant (2023) prove that the NPG algorithm converges to the optimal policy in the finite-state average-reward setting at rate $O(1/T)$. To do so, the paper assumes that there exists some uniform lower bound $\Delta$ on the relative state probability of an arbitrary policy and of the optimal policy, in stationarity:
@@
These proof approaches do not straightforwardly generalize to our infinite-state average-reward setting.
@@ 4.2. Proof sketch
Delving deeper into the Even-Dar et al. (2009) approach, their key insight is that one can think of the MDP optimization problem as consisting of many instances of the expert advice problem (Cesa-Bianchi et al., 1997), and think of the NPG algorithm as running a separate instance of the weighted majority algorithm for each state $s$, where the reward function at time step $k$ is $Q_{k}(s,a)$.
@@
Thus, to bound the convergence rate of the NPG algorithm, it is crucial to bound the relative value function $V_{\pi}(s)$, and thereby bound the state-action relative value $Q_{k}(s,a)$. In the finite-state setting, Even-Dar et al. (2009) use their bounded-mixing time assumption to prove a universal bound on $V_{\pi}(s)$ over all policies $\pi$, states $s$, and actions $a$, which allows the NPG convergence rate proof to be completed.
In our infinite-state setting, we also prove bounds on $V_{\pi}(s)$, but we prove policy- and state-dependent bounds. Specifically, we prove strong enough bounds to complete the proof in a similar fashion to Even-Dar et al. (2009), transferring bounds on the relative value function to bounds on the convergence rate. These relative-value bounds form the key novelty of this paper.
We now outline our lemmas, which primarily focus on bounding $V_{\pi}(s)$, towards our goal of proving Theorem 1, our convergence rate bound.
@@
We start with Lemma 5 by bounding how much worse $V_{k+1}(s)$ can be than $V_{k}(s)$, relative to the policy’s hitting time $\tau_{k}(s)$ from states $s$ to the special state $\vec{0}$. This lemma builds on Lemma 1, a standard monotonicity result for the NPG policy.
@@
Step 4: Main result, using the quadratic assumption   In Theorem 1, we specifically assume that the initial policy’s relative value function $V_{0}(s)$ grows at most quadratically relative to the reward $r(s,a)$, an assumption that we show in Section 6 is satisfied by the MaxWeight policy in the setting of queueing MDPs.
As a result, we now have a bound on the relative value function $V_{k}(s)$ for each NPG iterate $\pi_{k}$ which depends on the state $s$ – the bound is not uniform over all states. We specify this bound in Lemma 8. Correspondingly, we set our learning rate $\beta$ as a function of the state $s$, based on the bound on the relative value function that we are able to prove in that particular state. In states with more negative reward, our bound on $V_{\pi}(s)$ is weaker, so we use a slower learning rate, to improve convergence.
@@ 5. Proofs
    In Section 5.5, we prove our main result, Theorem 1, making use of our quadratic-relative-value assumption on our initial policy $\pi_{0}$.
@@ 5.1. Background lemmas
We start by restating results from the literature. Note that Lemmas 1 and 2 were each originally proven in a setting with fixed $\beta$, but the proofs still hold unchanged in our setting with variable $\beta_{s}$.
@@
Lemma 0 ().
@@
In (Murthy and Srikant, 2023, Lemma 2). We restate the proof in Appendix A for added clarity. ∎
Lemma 0 ().
Average reward increases monotonically: $J_{k+1}\geq J_{k}$.
Lemma 2 follows immediately from Lemma 1 and the Performance Difference Lemma, Lemma 3, which relates the difference in average reward between two policies to their relative value functions:
Lemma 0 ().
(Cao, 1999, (10)) For any pair of policies $\pi,\pi^{\prime}$,
@@
In particular, Lemma 3 holds in our infinite-state setting. The proof of Lemma 3 given by Cao (1999) requires only that the underlying Markov chains are ergodic, ensuring that the relevant quantities are well-defined.
@@ 5.2. Step 1: Bounding change in relative value under NPG update
First, we bound the amount by which the relative value function of an NPG iterate $V_{k+1}(s)$ can be lower (i.e. worse) than $V_{k}(s)$, the previous policy iterate, relative to $\tau_{k+1}(s)$, the time to hit $\vec{0}$ from state $s$ under policy $\pi_{k+1}$:
Lemma 0 ().
@@
$V_{k}^{\tau_{k+1}(s)}(s)$ $=\mathbb{E}\left[\sum\nolimits_{i=0}^{\tau_{k+1}(s)}\left(r(s_{k+1}(i),a_{k+1}(i))-J_{k}\right)\right]=V_{k+1}(s)+\tau_{k+1}(s)(J_{k+1}-J_{k})$
Applying the fact that $V_{k}^{\tau_{k+1}(s)}(s)\geq V_{k}(0)$, we find that
$V_{k+1}(s)+\tau_{k+1}(s)(J_{k+1}-J_{k})$ $=V_{k}^{\tau_{k+1}(s)}(s)\geq V_{k}(0)\implies V_{k+1}(s)\geq V_{k}(0)-\tau_{k+1}(s)(J_{k+1}-J_{k}).\qed$
5.3. Step 2: Bounding the time to hit the highest-reward state $\vec{0}$.
Now, we bound the time $\tau_{\pi}(s)$ for the system to hit the highest reward state $\vec{0}$, starting from any state $s$ and under any policy $\pi$. Later, we will apply this bound when the initial state $s$ has high reward $r_{\max}(s)$ and for policies $\pi$ with high average reward $J_{\pi}(s)$. In doing so, we will make key use of Assumption 2, our assumption of uniform connectedness of states with high reward $r_{\max}(s)$.
Lemma 0 ().
Let $y,z$ be two reward thresholds, $y>z$. For all states $s$ with $r_{\max}(s)\geq z$ and all policies $\pi$ such that $J_{\pi}\geq y$, the hitting time is bounded:
$\tau_{\pi}(s)\leq\tau^{bound}_{y,z}:=\frac{x_{z}(c_{\max}-y)}{p_{z}^{2}(y-z)}+x_{z}$
@@
Let $\mathbb{S}_{z}$ be the set of states $s$ such that $r_{\max}(s)\geq z$. For each state $s\in\mathbb{S}_{z}$, let $t_{\pi,z}(s)$ be the expected amount of time until policy $\pi$ next reaches a state $s^{\prime}\in\mathbb{S}_{z}$, starting in state $s$. Let $t_{\pi,z}^{\max}$ be the maximum over $s\in\mathbb{S}_{z}$ of $t_{\pi,z}(s)$. Let $s^{*}$ be the state in $\mathbb{S}_{z}$ which achieves this maximum. By Assumption 2, every state $s\in\mathbb{S}_{z}$ has at least $p_{z}$ probability of moving to state $s^{*}$ in the next $x_{z}$ steps, and therefore spending at most $x_{z}$ of the next $t_{\pi,z}^{\max}+x_{z}$ steps inside $\mathbb{S}_{z}$.
Let us use a renewal-reward argument, where the renewal moments are steps on which the system visits state $s^{*}$. Let us subdivide this renewal period into three segments:
  1. (1)
    Starting at $s^{*}$, the time until a state $s\in S_{z}$ is next visited.
  2. (2)
    Starting from a state $s\in S_{z}$, the next $x_{z}$ steps, during which there is at least a $p_{z}$ probability of reaching the state $s^{*}$.
  3. (3)
    Any time not in either of these two segments, which must be spent outside of $S_{z}$.
From these segments, we can upper-bound the fraction of time during the renewal period that the system spends in $\mathbb{S}_{z}$: The system can only be in $\mathbb{S}_{z}$ during Item 2, and the system can spend at most $\frac{1}{p_{z}}$ periods of length at most $x_{z}$ in that segment. The fraction of time spent in $\mathbb{S}_{z}$ is maximized if no time is spent in Item 3, resulting in the following bound:
$\mathbb{P}\{s_{t}\in\mathbb{S}_{z}\}\leq\frac{x_{z}/p_{z}}{t_{\pi,z}^{\max}+x_{z}/p_{z}}$
Recalling that the system accrues at most $c_{\max}$ reward inside of $\mathbb{S}_{z}$ and at most $z$ reward outside of $z$, we can conclude that the long-term reward of the policy $\pi$ is at most
(3) $J_{\pi}\leq c_{\max}\frac{x_{z}/p_{z}}{t_{\pi,z}^{\max}+x_{z}/p_{z}}+z\frac{t_{\pi,z}^{\max}}{t_{\pi,z}^{\max}+x_{z}/p_{z}}=\frac{x_{z}c_{\max}+p_{z}t_{\pi,z}^{\max}z}{x_{z}+p_{z}t_{\pi,z}^{\max}}$
However, we know that $J_{\pi}\geq y$. We can therefore use (3) to bound $t_{\pi,z}^{\max}$:
$y$ $\leq\frac{x_{z}c_{\max}+p_{z}t_{\pi,z}^{\max}z}{x_{z}+p_{z}t_{\pi,z}^{\max}}\implies t_{\pi,z}^{\max}\leq\frac{x_{z}(c_{\max}-y)}{p_{z}(y-z)}$
Now, we can bound the time until we reach $\vec{0}$, starting from $s$. Whenever the system is in a state in $\mathbb{S}_{z}$, there is a $p_{z}$ chance of reaching $\vec{0}$ in the next $x_{z}$ steps, and a $t_{\pi,z}^{\max}$ expected time until we next get the opportunity.
Let $p_{\pi,z}(s)$ be the probability that the system reaches $\vec{0}$ from $s$ in at most $x_{z}$ steps. Let $t^{\prime}_{\pi,z}(s)$ be the expected time until the system re-enters $\mathbb{S}_{z}$, conditional on not reaching $\vec{0}$ in at most $x_{z}$ steps. Note that by Markov’s inequality, $t^{\prime}_{\pi,z}(s)\leq\frac{t_{\pi,z}(s)}{1-p_{\pi,z}(s)}$. We can now start to bound $\tau_{\pi}(s)$ as follows:
$\tau_{\pi}(s)\leq p_{\pi,z}(s)x_{z}+(1-p_{\pi,z}(s))(t^{\prime}_{\pi,z}(s)+\tau_{\pi}(s^{\prime}))$
@@
$\tau_{\pi}(s)$ $\leq p_{\pi,z}(s)x_{z}+(1-p_{\pi,z}(s))\left(\frac{t_{\pi,z}(s)}{1-p_{\pi,z}(s)}+\tau_{\pi}(s^{\prime})\right)=x_{z}+t_{\pi,z}(s)+(1-p_{\pi,z}(s))(\tau_{\pi}(s^{\prime})-x_{z})$
In particular, letting $\tau_{\pi}^{\max}$ be the maximum over $s\in\mathbb{S}_{z}$ of $\tau_{\pi}(s)$, and letting $s_{\tau}^{*}$ be the state in which that maximum occurs, we have
$\tau_{\pi}^{\max}$ $\leq x_{z}+t_{\pi,z}(s_{\tau}^{*})+(1-p_{\pi,z}(s_{\tau}^{*}))(\tau_{\pi}^{\max}-x_{z})$
$\tau_{\pi}^{\max}$ $\leq x_{z}+\frac{t_{\pi,z}(s_{\tau}^{*})}{p_{\pi,z}(s_{\tau}^{*})}\leq x_{z}+\frac{t_{\pi,z}^{\max}}{p_{z}}=\frac{x_{z}(c_{\max}-y)}{p_{z}^{2}(y-z)}+x_{z}.\qed$

### v2

@@ 4. Results
We prove the first convergence result for the Natural Policy Gradient (NPG) algorithm in the setting of unbounded queues. In particular, we show that for any queueing model in the Generalized Switch with Static Environment (GSSE) class of models, the NPG algorithm rapidly converges to the scheduling policy with minimal mean queue length. The GSSE setting is a broad, natural family of queueing systems, which includes as special cases such important queueing systems as the $n\times n$ switch, the N-system, and the multiserver-job system.
To achieve this rapid convergence, we initialize the NPG algorithm with the MaxWeight scheduling policy, a well-studied policy that is known to achieve optimal stability region, but which does not achieve optimal or near-optimal mean queue length except in certain asymptotic regimes. We also select a specific learning rate parameterization $\beta_{s}$, whose properties are vital to our convergence result.
We demonstrate that after $T$ iterations, the $T$th NPG iterate will achieve a mean queue length within $O(1/\sqrt{T})$ of the optimal policy.
Theorem 1.
For any GSSE queueing model with the objective of minimizing mean queue length, the NPG algorithm with initial policy $\pi_{0}=MaxWeight$ and with learning rate parameterization given in Theorem 2 achieves the following mean queue length:
(4) $E[|q_{\pi_{T}}|_{1}]\leq E[|q_{\pi^{*}}|_{1}]+\frac{c_{*}}{\sqrt{T}},$
where $\pi_{T}$ is the $T$th iterate of the NPG algorithm, $\pi^{*}$ is the policy with minimal mean queue length, and $c_{*}$ is a constant depending on the specific GSSE model.
This result generalizes to the $\alpha$th moment of queue length.
Proof deferred to Section 5.8..
Beyond the GSSE setting, our approach holds for a more general family of infinite-state average-reward MDPs, including both queueing systems outside the GSSE family of models, as well as a wide variety of non-queueing systems with countably infinite state spaces. This is the first NPG convergence result for any general class of infinite-state average-reward MDPs.
We now state our result in full generality:
Theorem 2.
@@
Proof deferred to Section 5..
We first prove Theorem 2, our more general result, and then apply it to the GSSE queueing setting to prove Theorem 1.
In Section 4.1, we discuss the key challenges which we needed to overcome to prove Theorem 1 and Theorem 2. We discuss our approach in Section 4.2, which centers on bounding the growth rate of the relative value function $V_{k}(s)$ under the NPG iterates $\pi_{k}$.
@@ 4.1. Challenges
We seek to prove a convergence rate result for the NPG algorithm in a setting of unbounded queues, which is an infinite-state-space, average-reward setting. A natural approach would be to try to generalize finite-state average-reward results on the NPG algorithm to the infinite-reward setting. There are two relevant approaches to consider: the approach of Even-Dar et al. (Even-Dar et al., 2009), and of Murthy and Srikant (Murthy and Srikant, 2023). Unfortunately, each of these approaches makes crucial use of assumptions that are only plausible for MDPs with finite state spaces.
With respect to Even-Dar et al. (Even-Dar et al., 2009), combining that paper’s Algorithm 4 (MDP Experts) with its Algorithm 1 (Weighted Majority) results in exactly the Natural Policy Gradient algorithm, and the paper’s Theorem 4.1 proves that NPG converges to the optimal policy in the finite-state average-reward setting, with convergence rate $O(1/\sqrt{T})$. To prove this theorem, the paper assumes that there exists some finite mixing time $\tau$ for all policies $\pi$. Intuitively, this assumption states that for every policy and every initial system state, the system evolves to a state distribution becomes near the policy’s stationary distribution in at most $O(\tau)$ steps. In a finite-state MDP, this is a reasonable assumption: For reasonable policies, $\tau$ might be exponential in $|S|$ at worst. In contrast, in a queueing system with unbounded queue length, this assumption is plainly false. No policy can achieve a finite mixing time, much less all policies.
Murthy and Srikant (Murthy and Srikant, 2023) prove that the NPG algorithm converges to the optimal policy in the finite-state average-reward setting at rate $O(1/T)$. To do so, the paper assumes that there exists some uniform lower bound $\Delta$ on the relative state probability of an arbitrary policy and of the optimal policy, in stationarity:
@@
These proof approaches do not straightforwardly generalize to our setting of unbounded queues.
@@ 4.2. Proof sketch
Delving deeper into the Even-Dar et al. (Even-Dar et al., 2009) approach, their key insight is that one can think of the MDP optimization problem as consisting of many instances of the expert advice problem (Cesa-Bianchi et al., 1997), and think of the NPG algorithm as running a separate instance of the weighted majority algorithm for each state $s$, where the reward function at time step $k$ is $Q_{k}(s,a)$.
@@
Thus, to bound the convergence rate of the NPG algorithm, it is crucial to bound the relative value function $V_{\pi}(s)$, and thereby bound the state-action relative value $Q_{k}(s,a)$. In the finite-state setting, Even-Dar et al. (Even-Dar et al., 2009) use their bounded-mixing time assumption to prove a universal bound on $V_{\pi}(s)$ over all policies $\pi$, states $s$, and actions $a$, which allows the NPG convergence rate proof to be completed.
In our infinite-state setting, we also prove bounds on $V_{\pi}(s)$, but we prove policy- and state-dependent bounds. Specifically, we prove strong enough bounds to complete the proof in a similar fashion to Even-Dar et al. (Even-Dar et al., 2009), transferring bounds on the relative value function to bounds on the convergence rate. These relative-value bounds form the key novelty of this paper.
We now outline our lemmas, which primarily focus on bounding $V_{\pi}(s)$, towards our goal of proving Theorem 2, our generalized convergence rate bound, and ultimately proving Theorem 1, our convergence rate bound for GSSE queues specifically.
Step 0: Prove that each iterate $\pi_{k}$ is stable   It is vital for our result that at every NPG iteration $k$, the current policy $\pi_{k}$ is stable. This is not trivial, as there exist unstable policies $\pi$ within our space of eligible policies.
In Lemma 5, we inductively prove that each policy $\pi_{k}$ is stable by using the relative value function $V_{k-1}(s)$ of the previous policy $\pi_{k-1}$ as a Lyapunov function.
@@
We start with Lemma 6 by bounding how much worse $V_{k+1}(s)$ can be than $V_{k}(s)$, relative to the policy’s hitting time $\tau_{k}(s)$ from arbitrary states $s$ to the special state $\vec{0}$. This lemma builds on Lemma 1, a standard monotonicity result for the NPG policy.
@@
Step 4: Main general result, using the quadratic assumption   In Theorem 2, we specifically assume that the initial policy’s relative value function $V_{0}(s)$ grows at most quadratically relative to the reward $r(s,a)$, an assumption that we show in Lemma 11 is satisfied by the MaxWeight policy in the setting of queueing MDPs.
As a result, we now have a bound on the relative value function $V_{k}(s)$ for each NPG iterate $\pi_{k}$ which depends on the state $s$ – the bound is not uniform over all states. We specify this bound in Lemma 9. Correspondingly, we set our learning rate $\beta$ as a function of the state $s$, based on the bound on the relative value function that we are able to prove in that particular state. In states with longer queues (more negative reward), our bound on $V_{\pi}(s)$ is weaker, so we use a slower learning rate, to improve convergence.
@@
Step 5: Main queueing result
Finally, in Theorem 1, we prove that every queueing system in the GSSE family of queues satisfies the requirements for Theorem 1 to apply, and that the MaxWeight policy’s relative value function $V_{MW}(s)$ grows at most quadratically with respect to total queue length. As a result, we show that the NPG algorithm is guaranteed to achieve a $O(1/\sqrt{T})$ convergence rate for any queueing model in the GSSE family.
@@ 5. Proofs
    In Section 5.6, we prove our main result stated in full generality, Theorem 2, making use of our quadratic-relative-value assumption on our initial policy $\pi_{0}$.
  6. (6)
    In Section 5.7, we prove that the MaxWeight policy achieves a quadratically-growing relative value function for any queueing system in the Generalized Switch with Static Environment (GSSE) family of queues.
  7. (7)
    Finally, in Section 5.8, we prove Theorem 1, our main result stated specifically for GSSE queues.
@@ 5.1. Background lemmas
We start by restating results from the literature. Note that Lemmas 1 and 3 were each originally proven in a setting with fixed $\beta$, but the proofs still hold unchanged in our setting with variable $\beta_{s}$.
@@
Lemma 0.
NPG iterate policy $\pi_{k+1}$ improves upon policy $\pi_{k}$ relative to $\pi_{k}$’s relative-value function $Q_{k}$:
@@
This result is proven for the case of non-state-dependent learning rate $\beta$ in (Murthy and Srikant, 2023, Lemma 2). We demonstrate that the proof transfers to the state-dependent case in Appendix A. The proof holds for infinite MDPs, and does not require $\pi_{k+1}$ to be a stable policy. ∎
Next, we state the Performance Difference Lemma, Lemma 2, which relates the difference in average reward between two policies to their relative value functions:
Lemma 0.
Performance Difference Lemma: (Cao, 1999, (10)) For any pair of stable policies $\pi,\pi^{\prime}$,
@@
In particular, Lemma 2 holds in our infinite-state setting. The proof of Lemma 2 given by Cao (Cao, 1999) requires only that the underlying Markov chains are ergodic, ensuring that the relevant quantities are well-defined.
Lemma 0.
If NPG iterate policies $\pi_{k}$ and $\pi_{k+1}$ are stable, then average reward increases monotonically: $J_{k+1}\geq J_{k}$.
Proof.
Lemma 3 follows immediately from Lemma 1 and Lemma 2. ∎
@@
5.2. Step 0: Stability
We start by showing that each NPG iterate is stable, via induction over the iterations.
Lemma 0.
For any Markov chain satisfying Assumption 1, if the NPG algorithm is initialized with a stable policy $\pi_{0}$, then all NPG iterates $\pi_{k}$ are stable.
Proof.
We will proceed by induction. We must show that if an NPG iterate $\pi_{k}$ is stable, then the next NPG iterate $\pi_{k+1}$ is stable.
We will use the function $-V_{k}(s)$ as a Lyapunov function for the policy $\pi_{k+1}$, and apply the Foster-Lyapunov theorem (Foster, 1953). To show that $\pi_{k+1}$ is a stable policy with $-V_{k}(s)$ as our Lyapunov function, we must show that
  • $V_{k}(s)$ is upper bounded by a constant for all $s$,
  • There exists a constant $C$ such that for all $s$, the expected one-step update of $V_{k}(s)$ is lower bounded:
    $\forall s,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)\geq C,$
  • There exists a finite set $S$ and positive constant $\epsilon>0$ such that the expected one-step update of $V_{k}(s)$ is bounded above zero:
    $\forall s\not\in S,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)\geq\epsilon.$
Let’s start by upper bounding $V_{k}(s)$. Note that the maximum value of $V_{k}(s)$ must be attained by a state $s$ such that the maximum possible reward $r_{\max}(s)$ exceeds $J_{k}$, as can be seen by examining the Poisson equation (1). There are only finitely many such states, by Assumption 1(b), so $V_{k}(s)$ is upper bounded.
Next, let’s look at the expected value of $V_{k}(s)$, after a one-step transition according to policy $\pi_{k+1}$, and apply (2), the Poisson equation:
(7) $E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]=Q_{k}(s,\pi_{k+1})+J_{k}-r(s,a).$
Now, we apply Lemma 1, which states that when measured according to $\pi_{k}$’s relative value function $Q_{k}(s,a)$, policy $\pi_{k+1}$’s actions have average relative value at least that of $\pi_{k}$:
$\forall s,\,Q_{k}(s,\pi_{k+1})\geq Q_{k}(s,\pi_{k})=V_{k}(s).$
Note that the Lemma 1 does not require $\pi_{k+1}$ to be a stable policy, so we may use Lemma 1. Combining Lemma 1 with (7),
$E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]\geq V_{k}(s)+J_{k}-r(s,a).$
Let $S$ be the set of states $s$ with maximum reward $r_{\max}(s)\geq J_{k}-1$. Note that $S$ is a finite set, by Assumption 1(b).
We then have the following bounds on the one-step update of $V_{k}(s)$ under policy $\pi_{k+1}$:
$\forall s,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)$ $=J_{k}-r(s,a)\geq J_{k}-c_{\max}$
$\forall s\not\in S,\,E_{s^{\prime}\sim P_{k+1}(s)}[V_{k}(s^{\prime})]-V_{k}(s)$ $=J_{k}-r(s,a)\geq J_{k}-(J_{k}-1)=1.$
Thus, all conditions on $V_{k}(s)$ are satisfied, and by the Foster-Lyapunov theorem, policy $\pi_{k+1}$ is stable. By induction, all NPG iterates are stable. ∎
@@ 5.3. Step 1: Bounding change in relative value under NPG update
Next, we bound the amount by which the relative value function of an NPG iterate $V_{k+1}(s)$ can be lower (i.e. worse) than $V_{k}(s)$, the previous policy iterate, relative to $\tau_{k+1}(s)$, the time to hit $\vec{0}$ from state $s$ under policy $\pi_{k+1}$:
Lemma 0.
@@
$V_{k}^{\tau_{k+1}(s)}(s)$ $=\mathbb{E}\left[\sum\nolimits_{i=0}^{\tau_{k+1}(s)}\left(r(s_{k+1}(i),a_{k+1}(i))-J_{k}\right)\right]$
$=V_{k+1}(s)+\tau_{k+1}(s)(J_{k+1}-J_{k}).$
Applying the fact that $V_{k}^{\tau_{k+1}(s)}(s)\geq V_{k}(s)$, we find that
$V_{k+1}(s)+\tau_{k+1}(s)(J_{k+1}-J_{k})$ $=V_{k}^{\tau_{k+1}(s)}(s)\geq V_{k}(s)$
$V_{k+1}(s)$ $\geq V_{k}(s)-\tau_{k+1}(s)(J_{k+1}-J_{k}).$
5.4. Step 2: Bounding the time to hit the empty-queue state $\vec{0}$.
Now, we bound the time $\tau_{\pi}(s)$ for the system to hit the empty queue state $\vec{0}$, or more generally the highest-reward state, starting from any state $s$ and under any policy $\pi$. Later, we will apply this bound when the initial state $s$ has high reward $r_{\max}(s)$ and for policies $\pi$ with high average reward $J_{\pi}$. In proving our bound, we make key use of Assumption 2, our assumption of uniform connectedness of states with high reward $r_{\max}(s)$, e.g. short queues.
Lemma 0.
Consider an MDP satisfying Assumptions 2 and 1. Let $y,z$ be two reward thresholds, $y>z$. For all states $s$ with $r_{\max}(s)\geq z$ and all stable policies $\pi$ such that $J_{\pi}\geq y$, the hitting time is bounded:
$\tau_{\pi}(s)\leq\tau^{bound}_{y,z}:=\frac{x_{z}(c_{\max}-y)}{p_{z}^{2}(y-z)}+x_{z},$
@@
Let $s$ be a generic state such that $r_{\max}(s)\geq z$. As $\pi$ is stable, it returns to $s$ infinitely often. Let us consider the renewal cycle with renewal point in state $s$.
Let $\mathbb{S}_{z}$ be the set of states $s^{\prime}$ with $r_{\max}(s^{\prime})\geq z$.
We will start by upper bounding the fraction of each cycle that is spent in states $s_{t}\not\in\mathbb{S}_{z}$. This will be used to construct an upper bound on the hitting time to reach the identified state $\vec{0}$ (e.g. the empty queue).
Let $T_{in}^{s}$ denote the expected time per $s$-cycle that the system spends in states $s_{t}\in\mathbb{S}_{z}$, and let $T_{out}^{s}$ denote the expected time per cycle with $s_{t}\not\in\mathbb{S}_{z}$. We can bound the total expected reward per cycle as
$\mathbb{E}[\text{total reward}]\leq c_{\max}T^{s}_{in}+zT^{s}_{out}$
Knowing that the expected mean reward $J_{\pi}$ is at least $y$, we can apply renewal reward:
(8) $y$ $\leq J_{\pi}=\frac{\mathbb{E}[\text{total reward}]}{T_{in}^{s}+T^{s}_{out}}\leq\frac{c_{\max}T^{s}_{in}+zT^{s}_{out}}{T^{s}_{in}+T^{s}_{out}}\implies T^{s}_{out}\leq\frac{c_{\max}-y}{y-z}T^{s}_{in}$
Next, we’ll directly prove an upper bound on $T^{s}_{in}$, which in turn upper bounds $T^{s}_{out}$. By Assumption 2, starting on each time step $t$ on which $s_{t}\in\mathbb{S}_{z}$, there is at least a probability $p_{z}$ that in the interval consisting of the next $x_{z}$ steps, the system reaches the renewal state $s$. There can be at most $\frac{1}{p_{z}}$ such intervals in expectation per renewal period, and all time steps on which $s_{t}\in\mathbb{S}_{z}$ must be contained in these intervals. As a result, $T^{s}_{in}\leq\frac{x_{z}}{p_{z}}$. Applying (8), we find that
$T^{s}_{out}\leq\frac{c_{\max}-y}{y-z}\frac{x_{z}}{p_{z}}$
Note that this bound on $T^{s}_{out}$ applies for any state $s\in\mathbb{S}_{z}$. Note also that $T^{s}_{out}$ is an upper bound on the expected time between visits to $\mathbb{S}_{z}$.
Let $p_{\pi,z}(s)$ be the probability that the system reaches $\vec{0}$ from a generic state $s\in\mathbb{S}_{z}$ in at most $x_{z}$ steps. Let $\{T^{s}_{out}\}^{\prime}$ be the expected time, starting from state $s$, until the system re-enters $\mathbb{S}_{z}$, conditional on not reaching $\vec{0}$ in the first $x_{z}$ steps. Note that
${T^{s}_{out}}^{\prime}\leq\frac{T^{s}_{out}}{1-p_{\pi,z}(s)},$
because the probability of the event of not reaching $\vec{0}$ in at most $x_{z}$ steps is $1-p_{\pi,z}(s)$. We can now start to bound $\tau_{\pi}(s)$ as follows, based on whether or not we hit $\vec{0}$ on the current visit to $\mathbb{S}_{z}$:
$\tau_{\pi}(s)\leq p_{\pi,z}(s)x_{z}+(1-p_{\pi,z}(s))({T^{s}_{out}}^{\prime}+\tau_{\pi}(s^{\prime}))$
@@
$\tau_{\pi}(s)$ $\leq p_{\pi,z}(s)x_{z}+(1-p_{\pi,z}(s))\left(\frac{T^{s}_{out}}{1-p_{\pi,z}(s)}+\tau_{\pi}(s^{\prime})\right)$
$=x_{z}+t_{\pi,z}(s)+(1-p_{\pi,z}(s))(\tau_{\pi}(s^{\prime})-x_{z})$
In particular, letting $\tau_{\pi}^{\max}$ be the maximum over $s\in\mathbb{S}_{z}$ of $\tau_{\pi}(s)$, and letting $s^{*}$ be the state in which that maximum occurs, we have
$\tau_{\pi}^{\max}$ $\leq x_{z}+T^{s^{*}}_{out}+(1-p_{\pi,z}(s^{*}))(\tau_{\pi}^{\max}-x_{z})$
$\tau_{\pi}^{\max}$ $\leq x_{z}+\frac{T^{s^{*}}_{out}}{p_{\pi,z}(s^{*})}\leq\frac{x_{z}(c_{\max}-y)}{p_{z}^{2}(y-z)}+x_{z}.$